OpenStudy (anonymous):

Solve the following equation sin2 x = -2cos x + 2

OpenStudy (freckles):

that is sin(2x) right?

OpenStudy (freckles):

on the left hand side ?

OpenStudy (anonymous):

its sin^2(x)

OpenStudy (freckles):

\[\sin^2(x)=-2\cos(x)+2 \text{ write \in terms of just } \cos \\ 1-\cos^2(x)=-2\cos(x)+2\]

OpenStudy (freckles):

this is a quadratic in terms of the function cos(x)

OpenStudy (freckles):

put everything on one side

OpenStudy (freckles):

\[0=+\cos^2(x)-2\cos(x)+2-1 \\ 0=\cos^2(x)-2\cos(x)+1\]

OpenStudy (freckles):

you can factor the right hand side

OpenStudy (freckles):

do you know how to factor u^2-2u+1?

OpenStudy (anonymous):

nope. no idea how to do any of this

OpenStudy (freckles):

hint (a-b)^2=a^2-2ab+b^2 so u^2-2u+1 =?

OpenStudy (freckles):

if you really forgot how to factor quadratics you could use the quadratic formula

OpenStudy (freckles):

if \[a \cdot u^2+b \cdot u+c=0 \\ \text{ then } u=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (freckles):

can you identify from your equation what a,b, and c are?

OpenStudy (anonymous):

are b and c 2?

OpenStudy (freckles):

you have \[(\cos(x))^2-2(\cos(x))+1=0 \\ \text{ you are comparing it to } \\ a(u)^2+b(u)+c=0\]

OpenStudy (freckles):

what is the coefficient in front of (cos(x))^2 what is the coefficient in front of the (cos(x)) what is the constant term?

OpenStudy (anonymous):

-2 1 0

OpenStudy (freckles):

hmm so that means you are looking at a totally different equation like -2cos^2(x)+cos(x)+0=0

OpenStudy (freckles):